OpenStudy (nakia_perez152):

The vertex of a parabola is (-1.5, -12.5), and its y-intercept is (0, -8). The x-intercepts of the parabola are a. (-2,0) b. (-4,0) c. (2,0) d. (4,0) and a. (-1,0) b. (-2,0) c. (1,0) d. (2,0)

OpenStudy (zzr0ck3r):

Vertex for is \(y=a(x-h)^2+k\) where \((h,k)\) is the vertex. So \(y=a(x+1.5)^2-12.5\) When we put in \(0\) we get \(-8\) so we use this to solve for \(a\). \(-8=a(0+1.5)-12.5\implies 4.5=1.5a\implies a = \dfrac{4.5}{1.5}=3\) So \(y=3(x+1.5)^2-12.5\) Now set equal to \(0\) and solve. \(3(x+1.5)^2-12.5=0\implies x+1.5=\pm\sqrt{\dfrac{12.5}{3}} \implies x=\pm\sqrt{\dfrac{12.5}{3}}-1.5\)